# VideoSAVi: Self-Aligned Video Language Models without Human Supervision

### v2

@@ 3.4 Preference Pair Creation and Optimization
The initial (dispreferred) response $a_{0}$ and the revised (preferred) response $a_{1}$ form a preference pair for DPO. These pairs are then used to train the model via DPO with the following loss function:
@@
where $\pi_{\theta}$ is the model policy being trained, $\pi_{\text{ref}}$ is a reference policy derived from the initial model, and $\beta$ is a scaling hyperparameter. This optimization process increases the likelihood that the model assigns higher probability to the refined, more accurate responses ($a_{1}$) over the initial, flawed responses ($a_{0}$).
Through repeated iterations of the entire pipeline, initial response generation, self-critiquing, refinement, and preference optimization, VideoSAVi progressively improves its ability to generate accurate responses and to identify and correct its own reasoning errors. This iterative process creates a self-improving cycle that enhances both spatial and temporal reasoning capabilities without requiring external supervision.
Model TempCompass PerceptionTest NeXTQA MVBench EgoSchema LongVideoBench (1) Baseline Models InternVL2.5 (Chen et al., 2024) 68.3 62.2 77.0 69.8 52.0 57.8 + SFT 68.5 63.0 77.5 70.2 53.0 58.0 + SFT+ 68.7 64.5 78.3 71.6 54.5 58.1 + Hound-DPO (Zhang et al., 2024b) 66.8 61.0 74.8 64.2 48.5 54.3 + TPO (Li et al., 2025) 68.2 62.0 77.2 68.8 52.8 58.1 (2) State-of-the-Art Models VideoLLaMA2 (Cheng et al., 2024) 43.4 51.4 - 54.6 51.7 - Kangaroo (Liu et al., 2024a) - - - 61.0 - 54.8 LLaVA-NeXT-Video (Zhang et al., 2024c) 53.0 48.8 53.5 53.1 - 49.1 LLaVA-NeXT-Interleave (Li et al., 2024b) 54.1 51.2 67.0 46.5 51.0 44.8 Qwen2-VL (Wang et al., 2024b) 68.9 62.3 75.7 64.9 59.2 55.6 LLaVA-OneVision (Li et al., 2024a) 64.5 57.1 79.3 56.7 64.0 56.3 LLaVA-Video (Zhang et al., 2024d) 66.4 67.9 74.2 58.6 57.6 58.2 (3) Preference-Optimized Models LLaVA-Hound-DPO (Zhang et al., 2024b) 55.5 45.1 61.6 36.6 36.1 36.7 i-SRT (Ahn et al., 2024) 56.0 47.0 63.0 36.3 46.2 38.2 LLaVA-Video-TPO (Li et al., 2025) 66.6 66.3 77.8 56.7 58.0 58.3 VideoSAVi 69.1+0.8 66.1+3.9 80.6+3.6 74.0+4.2 58.8+6.8 59.8+2.0
Table 1: Comprehensive evaluation of VideoSAVi against leading video understanding models. Best scores are in bold, and second-best scores are underlined. For VideoSAVi, performance improvements over InternVL 2.5 are shown as subscripts. All results except those marked with † are reproduced using LMMs-Eval (Zhang et al., 2024a).
@@ 4 Experiments and Evaluations
For training data, we sample a diverse collection of 4,000 videos: 2,000 from Star (Wu et al., 2021), and 1,000 each from Vidal (Zhu et al., 2024) and WebVid (Bain et al., 2021). Our self-supervised pipeline yields a comprehensive dataset of 24,000 preference pairs (3 pairs each for spatial and temporal reasoning). VideoSAVi uses state-of-the-art InternVL2.5 (Chen et al., 2024) as its backbone and is optimized using the SWIFT (Zhao et al., 2024) framework. Training and evaluation are conducted on two NVIDIA L40S GPUs (48GB) with maximum frames set to 32 (to prevent CUDA OOM). We employ LoRA (Hu et al., 2021) with $\alpha=\beta=8$ and DPO scaling factor $\beta=0.1$. The model converges in one epoch, requiring 12 hours of training time. The evaluation uses LMMs-Eval (Zhang et al., 2024a) for a fair comparison to prior work. We perform four iterations of self-training. The appendix includes additional experiments on: iteration-wise results (§A.1), preference data composition (§A.2), training data composition (§A.3), DPO training dynamics (§A.4), qualitative analysis (§A.6), dataset samples (§A.7), critique examples (§A.8), and prompt design (§A.9).
@@
We evaluate on general video understanding benchmarks: MVBench (multi-task reasoning) (Li et al., 2024c), PerceptionTest (visual perception) (Patraucean et al., 2024), TempCompass (temporal understanding) (Liu et al., 2024b), and NeXTQA (compositional reasoning) (Xiao et al., 2021). For long-form evaluation, we use EgoSchema (egocentric 3-min long videos) (Mangalam et al., 2023) and LongVideoBench (hour-long videos) (Wu et al., 2024).
@@ 4.1 Results
On baseline models, VideoSAVi outperforms the foundation InternVL2.5 (Chen et al., 2024) by substantial margins across all benchmarks: +0.8% on TempCompass, +3.9% on PerceptionTest, +3.6% on NeXTQA, +4.2% on MVBench, +6.8% on EgoSchema, and +2.0% on LongVideoBench. Fine-tuning on VInstruct (Maaz et al., 2023) (i.e., SFT) and our aligned SFT+ with preferred responses show incremental improvements, but neither approach matches our model’s generalization capabilities. Critically, VideoSAVi overcomes the limitations of Hound-DPO (Zhang et al., 2024b), which demonstrates that purely text-based ranking of preferences is fundamentally inadequate for video understanding, leading to severe performance degradation on MVBench (-5.6%) and EgoSchema (-3.5%).
VideoSAVi sets new state-of-the-art on key benchmarks through self-alignment.
Our approach surpasses Qwen2-VL (Wang et al., 2024b) by +9.1% on MVBench achieving a remarkable 74.0% accuracy, setting a new state-of-the-art. On NeXTQA, VideoSAVi outperforms the previous best, LLaVA-OneVision (Li et al., 2024a), by +1.3 percentage points. While LLaVA-Video (Zhang et al., 2024d) maintains an edge on PerceptionTest (67.9% vs. 66.1%), and LLaVA-OneVision leads on EgoSchema (64.0% vs. 58.8%), VideoSAVi demonstrates superior overall generalization across the diverse set of benchmarks.
@@
Previous preference-optimized approaches show inconsistent performance across benchmarks - LLaVA-Hound-DPO (Zhang et al., 2024b) and i-SRT (Ahn et al., 2024) suffer large performance drops, while TPO (Li et al., 2025) fails to improve temporal understanding despite its specific focus. In contrast, VideoSAVi delivers consistent improvements, outperforming TPO by +1.5 percentage points on LongVideoBench and reaching 59.8%. These results fundamentally challenge the notion that preference optimization inherently sacrifices consistency. Our self-alignment approach demonstrates robust performance across all dimensions of video understanding, maintaining strong capabilities in both temporal reasoning and spatial comprehension where previous methods show significant variability.
@@ 4.2 Ablation Studies
Factual Inaccuracy 45 29 20 12
Temporal Ordering 38 22 15 8
Spatial Relationship 42 31 18 9
Object Hallucination 29 18 11 6
Causal Reasoning 25 14 6 4
Detail Omission 21 16 4 2
Total 200 130 74 41
@@
Training Dist. 76.1 73.8 72.5
Cross-Question 72.5-${}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-3.6}}}$ 66.0-${}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-7.8}}}$ 59.7-${}_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-12.8}}}$
Cross-Video 71.8-${}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-4.3}}}$ 65.2-${}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-8.6}}}$ 58.5-${}_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-14.0}}}$
Full OOD 70.2-${}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-5.9}}}$ 63.0-${}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-10.8}}}$ 55.8-${}_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-16.7}}}$
Adversarial 69.4-${}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-6.7}}}$ 61.5-${}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-12.3}}}$ 53.0-${}_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-19.5}}}$
Compositional 70.7-${}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-5.4}}}$ 62.2-${}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-11.6}}}$ 54.2-${}_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-18.3}}}$
Avg. Gen. Gap -5.2 -10.2 -16.3
@@
To address concerns about the potential of self-reinforcement of errors in our self-critique approach, we conduct an error propagation analysis across four iterations of VideoSAVi (Table 2). We select 200 video-question pairs generated by iteration 1 of our model and have each subsequent iteration model (2, 3, and 4) independently answer these same questions without access to previous responses. GPT-4o (Hurst et al., 2024) evaluates all responses using the prompt provided in Figure 12. This method tests whether the model experiences “self-delusion,” wherein a self-improving method might reinforce its own incorrect beliefs rather than genuinely improving. The results demonstrate a consistent and substantial error reduction across all categories, with total errors decreasing by 79.5% from iteration 1 (200 errors) to iteration 4 (41 errors). This error reduction directly correlates with our iteration-wise performance improvements (§A.1), confirming that VideoSAVi genuinely learns from and corrects its reasoning flaws rather than reinforcing them.
@@
Table 3 evaluates VideoSAVi’s ability to generalize beyond its training distribution. We design a comprehensive test bed (detailed methodology in Appendix §A.5) with increasingly challenging conditions: novel question formulations (Cross-Question), unseen videos (Cross-Video), new domains (Full OOD), deliberately challenging inputs (Adversarial), and cases requiring integrative reasoning (Compositional). VideoSAVi demonstrates exceptional robustness with an average generalization gap of only -5.2%. This resilience is particularly evident in the most challenging scenarios, where VideoSAVi maintains 70.7% accuracy on compositional questions and 69.4% on adversarial examples. In contrast, TPO (Li et al., 2025) shows a more substantial reduction (-10.2% average gap), particularly struggling with compositional reasoning (-11.6%), while Hound-DPO (H-DPO) (Zhang et al., 2024b) shows the largest generalization gap (-16.3%), with performance dropping by nearly 20% on adversarial examples. The small performance variance across conditions confirms that our self-alignment approach develops genuine reasoning capabilities rather than memorizing patterns from training data.
Impact of Preference Data Size.
(a) TempCompass
(b) PerceptionTest
(c) NeXTQA
(d) MVBench
(e) EgoSchema
(f) LongVideoBench
Figure 2: Impact of preference data size on model performance.
Figure 2 demonstrates how VideoSAVi’s self-aligning pipeline consistently improves performance as the percentage of preference pairs (from final iteration 4 of self-training) increases across all benchmarks. Unlike TPO (Li et al., 2025), which shows minimal improvement or plateaus (particularly on EgoSchema), and Hound-DPO (Zhang et al., 2024b), which exhibits significant performance degradation after initial gains, VideoSAVi maintains consistent improvements throughout training. This robustness stems from our self-critiquing mechanism that targets specific hard reasoning cases, effectively identifying subtle temporal inconsistencies and spatial reasoning errors. The steep improvement curve on EgoSchema (+6.8 points) highlights how our approach excels at complex ego-centric understanding tasks, while the consistent gains on LongVideoBench demonstrate effective temporal alignment even when competing methods struggle with longer-form content.
Human Evaluation.
To validate the quality of self-generated preference pairs, we conduct a rigorous human evaluation with 6 external evaluators across 516 videos (261 temporal and 255 spatial). Evaluators are shown videos along model-generated questions and preference pairs without revealing the reasoning category. A pair is considered correct if and only if: (1) the question is answerable from video content, (2) the preferred response is accurate, and (3) the dispreferred response contains clear errors. This strict protocol reveals strong performance for both spatial relationships (71.3%) and temporal ordering (67.1%), with an overall quality of 69.2%. These results are particularly notable for preference learning in video understanding. Moreover, since the fully self-supervised pipeline operates without human intervention or external models, these findings confirm that our self-critiquing mechanism effectively identifies genuine reasoning errors rather than arbitrary preferences. This underscores that self-alignment can generate high-quality training signals for complex video understanding tasks.
Model-Agnostic Improvements.
Model TempCompass PerceptionTest NeXTQA MVBench EgoSchema LongVideoBench VideoLLaVA (Lin et al., 2024) 34.3 42.6 58.4 34.1 18.8 39.1 + VideoSAVi 36.8 +2.5 45.9 +3.3 63.2 +4.8 38.7 +4.6 23.5 +4.7 41.8 +2.7 LLaVA-NeXT-Interleave (Li et al., 2024b) 53.2 51.0 67.3 46.5 51.0 44.8 + VideoSAVi 55.6 +2.4 54.4 +3.4 71.2 +3.9 50.3 +3.8 54.6 +3.6 47.9 +3.1 LLaVA-OneVision (Li et al., 2024a) 64.1 57.5 79.3 56.7 64.0 56.3 + VideoSAVi 66.3 +2.2 60.8 +3.3 80.9 +1.6 59.9 +3.2 65.7 +1.7 58.7 +2.4 Qwen2-VL (Wang et al., 2024b) 68.9 62.1 75.6 64.9 59.2 55.6 +VideoSAVi 68.4 -0.5 65.4 +3.3 78.8 +3.2 68.3 +3.4 62.8 +3.6 58.1 +2.5
Table 4: Model-agnostic improvements from VideoSAVi.
Table 4 demonstrates the broad applicability of our approach across diverse architectures. VideoSAVi consistently enhances performance, delivering substantial improvements on nearly all benchmarks, regardless of the base model. The gains are particularly notable for VideoLLaVA (Lin et al., 2024), with improvements of +4.8 points on NeXTQA, +4.7 points on EgoSchema, and +4.6 on MVBench, showing that models with greater headroom benefit most from our preference optimization. For mid-tier models like LLaVA-NeXT-Interleave (Li et al., 2024b), VideoSAVi provides well-balanced improvements (+3.9 on NeXTQA, +3.8 on MVBench, +3.6 on EgoSchema), strengthening both spatial and temporal reasoning capabilities. Notably, VideoSAVi also enhances top-performing models, improving LLaVA-OneVision (Li et al., 2024a) by +1.6 on NeXTQA, reaching an impressive 80.9% and boosting Qwen2-VL (Wang et al., 2024b) across five benchmarks. These consistent cross-architecture improvements (an average gain of +3.0% across all models and benchmarks) confirm that our self-aligning preference learning framework addresses fundamental challenges in video understanding rather than exploiting model-specific characteristics.
Externally Trained Critics vs. Self-Critiquing.
[…]
Parameter Scaling vs. Preference Learning.

### v3

@@ 3.4 Preference Pair Creation and Optimization
The initial (dispreferred) response $a_{0}$ and the revised (preferred) response $a_{1}$ form a preference pair for DPO. These pairs are then used to train the model with the following loss function:
@@
where $\pi_{\theta}$ is the model policy being trained, $\pi_{\text{ref}}$ is a reference policy derived from the initial model, and $\beta$ is a scaling hyperparameter. This objective encourages the model to assign higher probability to refined, more accurate responses ($a_{1}$) over the initial, flawed ones ($a_{0}$).
Through repeated iterations of the entire pipeline (i.e., initial response generation, self-critiquing, refinement, and preference optimization), VideoSAVi progressively improves its ability to generate accurate responses and to identify and correct its own reasoning errors. This iterative process creates a self-improving cycle that enhances both spatial and temporal reasoning capabilities without requiring external supervision.
Model TempCompass PerceptionTest NeXTQA MVBench EgoSchema LongVideoBench (1) Baseline Models InternVL2.5 (Chen et al., 2024) 68.3 62.2 77.0 69.8 52.0 57.8 + SFT 68.5 63.0 77.5 70.2 53.0 58.0 + SFT+ 68.7 64.5 78.3 71.6 54.5 58.1 + Hound-DPO (Zhang et al., 2024b) 66.8 61.0 74.8 64.2 48.5 54.3 + TPO (Li et al., 2025b) 68.2 62.0 77.2 68.8 52.8 58.1 (2) State-of-the-Art Models VideoLLaMA2 (Cheng et al., 2024) 43.4 51.4 - 54.6 51.7 - Kangaroo (Liu et al., 2024a) - - - 61.0 - 54.8 LLaVA-NeXT-Video (Zhang et al., 2024c) 53.0 48.8 53.5 53.1 - 49.1 LLaVA-NeXT-Interleave (Li et al., 2024b) 54.1 51.2 67.0 46.5 51.0 44.8 Qwen2-VL (Wang et al., 2024b) 68.9 62.3 75.7 64.9 59.2 55.6 LLaVA-OneVision (Li et al., 2024a) 64.5 57.1 79.3 56.7 64.0 56.3 LLaVA-Video (Zhang et al., 2024d) 66.4 67.9 74.2 58.6 57.6 58.2 (3) Preference-Optimized Models LLaVA-Hound-DPO (Zhang et al., 2024b) 55.5 45.1 61.6 36.6 36.1 36.7 i-SRT (Ahn et al., 2024) 56.0 47.0 63.0 36.3 46.2 38.2 LLaVA-Video-TPO (Li et al., 2025b) 66.6 66.3 77.8 56.7 58.0 58.3 VideoSAVi 69.1+0.8 66.1+3.9 80.6+3.6 74.0+4.2 58.8+6.8 59.8+2.0
Table 1: Comprehensive evaluation of VideoSAVi against leading video understanding models. Best scores are in bold, and second-best scores are underlined. For VideoSAVi, performance improvements over InternVL 2.5 are shown as subscripts (reported as absolute gains in percentage points). All results except those marked with † are reproduced using LMMs-Eval (Zhang et al., 2024a).
@@ 4 Experiments and Evaluations
For training data, we sample a diverse collection of 4,000 videos from three sources: 2,000 from Star (Wu et al., 2021), and 1,000 each from Vidal (Zhu et al., 2024) and WebVid (Bain et al., 2021). Our self-supervised pipeline yields a comprehensive dataset of 24,000 preference pairs (6 pairs per video: 3 spatial reasoning + 3 temporal reasoning $\times$ 4,000 videos). To ensure fair evaluation, we decontaminate the training data by removing any videos that overlap with the evaluation benchmarks. VideoSAVi uses state-of-the-art InternVL2.5 (Chen et al., 2024) as its backbone and is optimized using the SWIFT (Zhao et al., 2024) framework. We employ LoRA (Hu et al., 2021) with $\alpha=8$ (and rank $r=8$), and DPO with temperature $\beta=0.1$. All training and evaluation are conducted on two NVIDIA L40S GPUs (48GB), with a maximum of 32 frames per video to avoid CUDA out-of-memory errors. The full training process consists of four iterations of self-training. Each iteration takes approximately 48 GPU hours (24 hours for generating preferences and 24 hours for DPO training), resulting in a total of $\sim$192 GPU hours across all iterations. For evaluation, we use LMMs-Eval (Zhang et al., 2024a) to ensure consistent comparison with prior work. Additional experiments and analyses are provided in the appendix, including: iteration-wise performance (§A.1), preference data composition (§A.2), training data composition (§A.3), DPO training dynamics (§A.4), qualitative examples (§A.6), dataset samples (§A.7), critique examples (§A.8), and prompt design (§A.9).
@@
We evaluate VideoSAVi on a range of general video understanding benchmarks, each targeting a specific reasoning skill: MVBench for multi-task reasoning (Li et al., 2024c), PerceptionTest for visual perception (Patraucean et al., 2024), TempCompass for temporal understanding (Liu et al., 2024b), and NeXTQA for compositional reasoning (Xiao et al., 2021). To assess performance on long-form video understanding, we use EgoSchema, which features 3-minute egocentric videos (Mangalam et al., 2023), and LongVideoBench, which includes hour-long videos across diverse tasks (Wu et al., 2024).
@@ 4.1 Results
Compared to its foundation model, InternVL2.5 (Chen et al., 2024), VideoSAVi shows significant improvements across all benchmarks: +0.8 percentage points on TempCompass, +3.9 on PerceptionTest, +3.6 on NeXTQA, +4.2 on MVBench, +6.8 on EgoSchema, and +2.0 on LongVideoBench. Supervised fine-tuning (SFT) on VInstruct (Maaz et al., 2023) and our aligned SFT+ with preferred responses show modest improvements, but neither matches our model’s generalization capabilities. Notably, VideoSAVi overcomes the limitations of Hound-DPO (Zhang et al., 2024b), which demonstrates that purely text-based ranking of preferences is fundamentally inadequate for video understanding, leading to severe performance degradation on MVBench (-5.6) and EgoSchema (-3.5).
VideoSAVi outperforms previous state-of-the-art on key benchmarks through self-alignment.
Our approach surpasses Qwen2-VL (Wang et al., 2024b) by +9.1 percentage points on MVBench achieving 74.0% accuracy, demonstrating significant improvements over existing methods. On NeXTQA, VideoSAVi outperforms the previous best, LLaVA-OneVision (Li et al., 2024a), by +1.3 percentage points. While LLaVA-Video (Zhang et al., 2024d) maintains an edge on PerceptionTest (67.9% vs. 66.1%), and LLaVA-OneVision leads on EgoSchema (64.0% vs. 58.8%), VideoSAVi demonstrates stronger generalization across the full set of benchmarks.
@@
Previous preference-optimized approaches show inconsistent performance across benchmarks. LLaVA-Hound-DPO (Zhang et al., 2024b) and i-SRT (Ahn et al., 2024) suffer large performance drops, and TPO (Li et al., 2025b) fails to improve temporal understanding despite its specific focus. In contrast, VideoSAVi delivers consistent improvements, outperforming TPO by +1.5 percentage points on LongVideoBench and reaching 59.8%. These findings challenge the assumption that preference optimization necessarily compromises consistency. Our self-alignment approach demonstrates robust performance across all dimensions of video understanding, maintaining strong capabilities in both temporal reasoning and spatial comprehension, where previous methods show significant variability.
@@ 4.2 Ablation Studies
Factual Inaccuracy 287 198 127 73
Temporal Ordering 243 162 108 52
Spatial Relationship 208 147 92 48
Object Hallucination 169 118 71 33
Causal Reasoning 132 91 58 27
Detail Omission 118 87 44 16
Total 1157 803 500 249
@@
Training Dist. 74.5 71.2 69.8
Cross-Question 70.2-$_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-4.3}}}$ $_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-8.7}}} 56.2-$$_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-13.6}}}$
Cross-Video 69.8-$_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-4.7}}}$ 61.8-$_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-9.4}}}$ 54.1-$_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-15.7}}}$
Full OOD 68.7-$_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-5.8}}}$ 59.4-$_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-11.8}}}$ 51.3-$_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-18.5}}}$
Adversarial 67.9-$_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-6.6}}}$ 57.8-$_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-13.4}}}$ 48.5-$_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-21.3}}}$
Compositional 68.3-$_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{-6.2}}}$ 58.6-$_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{-12.6}}}$ 49.2-$_{{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textbf{-20.6}}}$
Avg. Gen. Gap -6.4 -13.5 -17.9
@@
To address concerns about potential error reinforcement in our self-critique approach, we conduct an error propagation analysis across four iterations of VideoSAVi (Table 2). We use 1200 video-question pairs generated by iteration 1 and have each subsequent iteration (2, 3, and 4) independently answer the same questions without access to previous responses. GPT-4o (Hurst et al., 2024) evaluates all answers using the prompt provided in Figure 12 (see Appendix). This setup tests for the risk of “self-delusion,” where a self-improving method might reinforce its own incorrect beliefs instead of correcting them. The results show a consistent and substantial reduction in errors, with total mistakes decreasing by 78.4%, from 1157 in iteration 1 to 249 in iteration 4. This reduction closely aligns with our iteration-wise performance improvements (Appendix §A.1), confirming that VideoSAVi learns to correct its reasoning rather than reinforcing errors.
@@
Table 3 evaluates VideoSAVi’s ability to generalize beyond its training distribution. We design a comprehensive test bed (detailed methodology in Appendix §A.5) with increasingly challenging conditions: novel question formulations (Cross-Question), unseen videos (Cross-Video), new domains (Full OOD), deliberately challenging inputs (Adversarial), and cases requiring integrative reasoning (Compositional). VideoSAVi demonstrates exceptional robustness with an average generalization gap of only -6.4 percentage points. This resilience is particularly evident in the most challenging scenarios, in which VideoSAVi maintains 68.3% accuracy on compositional questions and 67.9% on adversarial examples. In contrast, TPO (Li et al., 2025b) shows a larger average gap of -13.5 percentage points, particularly struggling with adversarial reasoning (-13.4), while Hound-DPO (H-DPO) (Zhang et al., 2024b) shows the largest generalization gap (-17.9 percentage points), with performance dropping by nearly 21 percentage points on adversarial examples. The small performance variance across conditions confirms that our self-alignment approach develops genuine reasoning capabilities rather than memorizing patterns from training data.
@@
Externally Trained Critics vs. Self-Critiquing.
Benchmark GPT-4o
TempCompass 73.8
PerceptionTest 72.1
NeXTQA 88.1
MVBench 64.6
EgoSchema 66.2
LongVideoBench 66.7
Table 6: GPT-4o performance on video understanding benchmarks.
[…]
Parameter Scaling vs. Preference Learning.